OpenStudy (anonymous):

Use basic identities to simplify the expression cos^2 θ/sin^2 θ +csc θ sin θ

OpenStudy (anonymous):

@TuringTest can you help?

OpenStudy (anonymous):

\[\frac{ \cos^2(\theta) }{ \sin^2(\theta) }+cosec (\theta)\sin (\theta) \]\[=\cot^2(\theta) + \frac{ 1 }{ \sin \theta }\sin(\theta) = \cot^2(\theta) + 1 =?\]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

the first what?

OpenStudy (anonymous):

Oh, lol got confused for a second...

OpenStudy (anonymous):

What trig identities do you know? :-)

OpenStudy (anonymous):

Uhh not really any, they don't really make sense to me

OpenStudy (anonymous):

Okay so I take it this is homework, if you look through your class notes, make a list of the trig statements you've been given and I'll help you work it out